OpenStudy (anonymous):

What is the solution set for |3x – 24| 27?

OpenStudy (amistre64):

you missed a sign :)

OpenStudy (anonymous):

x= -1,-17

OpenStudy (anonymous):

\[\le\]

OpenStudy (anonymous):

I assume you left =

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

-27<= 3x -24<=27

OpenStudy (amistre64):

in this case it would be good to see that the innards have to either be equal to 27 or -27 and solve for both cases